# Two-Sample Testing with Block-Wise Missingness in Multi-Source Data

### v1

@@ Abstract
Multi-source and multi-modal datasets are increasingly common in scientific research, yet they often exhibit block-wise missingness, where entire data sources or modalities are systematically absent for subsets of subjects. This structured form of missingness presents significant challenges for statistical analysis, particularly for two-sample hypothesis testing. Standard approaches such as imputation or complete-case analysis can introduce bias or result in substantial information loss, especially when the missingness mechanism is not random. To address this methodological gap, we propose the Block-Pattern Enhanced Test (BPET), a general framework for two-sample testing that directly accounts for block-wise missingness without requiring imputation or deletion of observations. As a concrete instantiation, we develop the Block-wise Rank In Similarity graph Edge-count (BRISE) test, which extends rank-based similarity graph methods to settings with block-wise missing data. Under mild conditions, we establish that the null distribution of BRISE converges to a $\chi^{2}$ distribution. Simulation studies show that BRISE consistently controls the type I error rate and achieves good statistical power under a wide range of alternatives. Applications to two real-world datasets with block-wise missingness further demonstrate the practical utility of our method in identifying meaningful distributional differences.
@@ 1 Introduction
Multi-source and multi-modal datasets are increasingly common in scientific research, encompassing diverse data types collected across individuals, institutions, and time. Examples include the Alzheimer’s Disease Neuroimaging Initiative (ADNI) (Mueller et al., 2005a, b), which combines imaging, genetic, and clinical data, and electronic health records (EHRs) (Johnson et al., 2023) that integrate medication histories and clinical notes.
While such integration enables more powerful and comprehensive analyses, it also poses challenges related to data quality and structure. A key issue is block-wise missingness, where entire data sources are absent for certain individuals (Figure 1). This often stems from clinical priorities, resource constraints, or institutional protocols. For instance, in ADNI, site-level practices can produce systematic gaps across modalities (Xue and Qu, 2021). Similar patterns appear in partial or incomplete multi-view learning (Li et al., 2014; Xu et al., 2015; Zhao et al., 2016; Zhang et al., 2020; Zhou et al., 2023), such as bilingual document clustering with missing translations or audiovisual tasks where only one modality is available.
@@
These challenges are especially acute in the context of two-sample testing, a fundamental task in statistics that seeks to determine whether two groups are drawn from the same distribution. Block-wise missingness introduces structural differences in data availability across groups that can distort analysis and bias conclusions. Formally, let $X_{1},\ldots,X_{m}\overset{i.i.d}{\sim}F_{X}$ and $Y_{1},\ldots,Y_{n}\overset{i.i.d}{\sim}F_{Y}$ be two independent samples. The goal is to test
$H_{0}:F_{X}=F_{Y}\quad\text{versus}\quad H_{a}:F_{X}\neq F_{Y}.$
When the data are fully observed, many powerful nonparametric methods exist, including graph-based tests (Friedman and Rafsky, 1979; Schilling, 1986; Henze, 1988; Rosenbaum, 2005; Chen and Zhang, 2013; Chen and Friedman, 2017; Chen et al., 2018; Zhou and Chen, 2023), classification-based tests (Lopez-Paz and Oquab, 2017; Kim et al., 2021; Hediger et al., 2022), interpoint distance-based tests (Biswas and Ghosh, 2014; Li, 2018), and kernel-based tests (Gretton et al., 2006; Eric et al., 2007; Gretton et al., 2009, 2012; Song and Chen, 2024). These approaches exploit ranks, distances, or reproducing kernel Hilbert space (RKHS) embeddings to characterize distributional differences and are effective in high-dimensional or non-Euclidean settings where parametric models are not feasible.
However, nonparametric tests for missing data are sparse; naive implementation of existing methods leads to inflated type I error or unsatisfied power. Figure 2 summarizes existing strategies using imputation, complete-case, or available-case analysis. Specifically, imputation method fills in missing entries using model-based predictions (Toutenburg, 1990; Zhang, 2016) but relies on strong assumptions and can bias inference if those assumptions fail. Complete-case analysis discards incomplete samples, which reduces effective sample size (Little, 1992; Ross et al., 2020). Available-case analysis uses only variables required for a given comparison but can still suffer information loss and inefficiency (Little, 1992).
@@
Recent work has extended classical two-sample testing procedures to handle entrywise missingness, but these approaches are often conservative under block-wise missingness. For example, MMD-Miss (Zeng et al., 2024a) and the univariate rank-based test of Zeng et al. (2024b) adopted bounding strategies that embed incomplete observations and construct finite upper and lower bounds for the maximum mean discrepancy (MMD) statistic (Gretton et al., 2012) and the Wilcoxon–Mann–Whitney statistic, respectively. The null is rejected only when both bounds fall in the rejection region, ensuring type I error control despite missing entries. In addition, Aleksić and Milošević (2025) proposes a weighted modification of the energy test statistic that leverages all available data under missingness. These approaches perform well for random missing, but they do not target structured block-wise mechanisms.
To fill the gap, we propose the Block-Pattern Enhanced Test (BPET), a general framework for two-sample testing under block-wise missing data. BPET partitions the dataset by missingness patterns, applies pattern-aware procedures to each compatible pattern pair using shared observed sources, and aggregates the results into a global test statistic. This framework allows BPET to directly operate on partially observed data without imputation or case deletion. Further, in our proposed framework, many testing methods can be chosen as the test statistic; in this paper, we focus on a graph and rank-based instantiation on the Rank In Similarity graph Edge-count (RISE) test (Zhou and Chen, 2023). We call this method the Block-wise Rank In Similarity graph Edge-count (BRISE) test, which adapts rank-weighted similarity graphs to block-wise missingness. In our implementation, we adopt the permutation-based inference to construct the null distribution of the test statistics. Unlike the standard permutation, which shuffles all labels while fixing only the total group sizes, we introduce a pattern-wise permutation that restricts permutations within each pattern and we derive a sufficient condition for its validity that is strictly weaker than missing completely at random (MCAR). In theory, under mild regularity conditions, we establish that the BRISE statistics have asymptotic $\chi^{2}$ null distributions. We also show that standard permutation can fail when the groups have unequal marginal pattern distributions. In simulations, BRISE maintains type I error control and achieves higher power than existing competitors; in two empirical studies, BRISE detects meaningful distributional differences.
In summary, our contributions are threefold. First, we introduce BPET, a general framework that conducts two-sample testing directly on partially observed multi-source data without imputation or case deletion. Second, we propose a pattern-wise permutation scheme and addresses a failure mode of standard permutation. Third, we instantiate BPET via BRISE, develop its asymptotic theory, and demonstrate competitive empirical performance in simulations and real applications.
@@
The remainder of the paper is organized as follows. Section 2 presents BPET and the proposed tests. Section 3 develops the theoretical properties. Sections 4 and 5 report simulation results and real data analyses. Section 6 concludes, with proofs and additional results in the Supplementary Materials.
2 Method
In this section, we first introduce the general framework and procedure of BPET in Section 2.1, then its adaptation to the RISE method in Section 2.2, and the construction of the test statistics tailored for block-wise missing data in Section 2.3.
2.1 Framework of Block-Pattern Enhanced Test (BPET)
Stage 1: Pattern Partition.
Suppose there are $L$ sources, indexed by $l=1,\dots,L$, where each source is defined over a domain $\mathcal{S}^{(l)}$. These domains may vary across sources. For example, $\mathcal{S}^{(l)}$ could be a Euclidean space such as $\mathbb{R}^{d_{l}}$, or a non-Euclidean space such as a network or a functional data space. Many real-world datasets, such as those from ADNI (Mueller et al., 2005a) described in Section 5.2, exhibit this multi-source structure by integrating diverse types of information, including imaging, genetic, and clinical measurements.
To accommodate block-wise missingness, let $n_{P}$ denote the total number of distinct missingness patterns observed in the data. Each pattern is represented as
@@ 2.1 Framework of Block-Pattern Enhanced Test (BPET)
where $P_{\alpha}^{(l)}=1$ if source $l$ is observed and 0 otherwise. For instance, when $L=3$, Figure 1 shows that $\mathcal{P}_{1}=(1,1,1)$ corresponds to complete observation; $\mathcal{P}_{2}=(1,1,0)$, $\mathcal{P}_{3}=(1,0,1)$, $\mathcal{P}_{4}=(0,1,1)$ correspond to cases with two observed sources; and $\mathcal{P}_{5}=(1,0,0)$, $\mathcal{P}_{6}=(0,1,0)$, $\mathcal{P}_{7}=(0,0,1)$ correspond to cases with only one observed source. In general, the number of possible nonempty patterns is at most $2^{L}-1$.
To unify our notations on two-sample problems, we introduce $\{Z_{i}\}_{i=1}^{N}$ referring to all pooled observations, where $\{Z_{i}\}_{i=1}^{m}=\{X_{j}\}_{j=1}^{m}$, $\{Z_{i}\}_{i=m+1}^{N}=\{Y_{j}\}_{j=1}^{n}$, and $N=m+n$. Thus, each observation can be written as $Z_{i}=(Z_{i}^{(1)},\ldots,Z_{i}^{(L)})$, where $Z_{i}^{(l)}\in\mathcal{S}^{(l)}$ is the $l$-th source or modality if observed. Further, the missingness pattern of $Z_{i}$ is $S_{i}=(S_{i}^{(1)},\ldots,S_{i}^{(L)})\in\{0,1\}^{L}$. Then the set of observations following pattern $\mathcal{P}_{\alpha}$ is
@@
Based on the notations above, we define $\mathcal{X}^{(\alpha)}=\mathcal{Z}^{(\alpha)}\cap\{X_{i}\}_{i=1}^{m}$ and $\mathcal{Y}^{(\alpha)}=\mathcal{Z}^{(\alpha)}\cap\{Y_{i}\}_{i=1}^{n}$, so that $\mathcal{Z}^{(\alpha)}=\mathcal{X}^{(\alpha)}\cup\mathcal{Y}^{(\alpha)}$. Let $m_{\alpha}=|\mathcal{X}^{(\alpha)}|$, $n_{\alpha}=|\mathcal{Y}^{(\alpha)}|$, and $N_{\alpha}=|\mathcal{Z}^{(\alpha)}|$, where $|\cdot|$ denotes the cardinality of a set. These satisfy
@@
During this stage, we partition our observation from both samples based on their patterns, which facilitates the implementation of Stage 2.
@@
Given the partitioned data, we define a pattern pair as $(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})$ for $1\leq\alpha,\beta\leq n_{P}$. For observations $Z_{i}\in\mathcal{Z}^{(\alpha)}$ and $Z_{j}\in\mathcal{Z}^{(\beta)}$, if their patterns share at least one source (i.e., $\mathcal{P}_{\alpha}\mathcal{P}_{\beta}^{\top}>0$), we define their dissimilarity as
@@
where $\rho_{l}(\cdot,\cdot)$ is a source-specific dissimilarity measure and $\mathtt{Norm}(\cdot)$ is a normalization function. Notice that the dissimilarity between a pair of observations only uses the shared sources or modalities observed; we will discuss how to compute dissimilarity measure if no shared sources. There are many possible choices of the source-specific dissimilarity measure $\rho_{l}(\cdot,\cdot)$, which leads to different test-specific procedures. For Euclidean sources, one may take $\rho_{l}(Z_{i}^{(l)},Z_{j}^{(l)})=\|Z_{i}^{(l)}-Z_{j}^{(l)}\|_{2}^{2}$ and $\mathtt{Norm}(x)=\sqrt{x}$, which yields the Euclidean distance over the shared dimensions. The formulation in (1) is very flexible, which allows heterogeneous multi-modal data by choosing $\rho_{l}(\cdot,\cdot)$ as, e.g., graph diffusion distances for networks (Hammond et al., 2013), cosine similarities for embeddings (Kenter and De Rijke, 2015), or tailored measures for categorical data (Boriah et al., 2008). As an illustration, in Figure 1, given $Z_{i}\in\mathcal{Z}^{(3)}$ and $Z_{j}\in\mathcal{Z}^{(5)}$ with $\mathcal{P}_{3}=(1,0,1)$ and $\mathcal{P}_{5}=(1,0,0)$, the only shared source is $\mathcal{S}^{(1)}$, and thus their dissimilarity is computed using only the first source or modality, which reduces to $\rho(Z_{i},Z_{j})=\mathtt{Norm}\big{(}{\rho_{1}(Z_{i}^{(1)},Z_{j}^{(1)})}\big{)}$.
If $\mathcal{P}_{\alpha}\mathcal{P}_{\beta}^{\top}=0$, there are no shared sources, and direct computation of dissimilarity is impossible. We adopt a simple solution by assigning zero values to such pairs. While this takes the risk to discard potentially useful information, empirically it performs well in both simulation studies and real data analysis (Sections 4 and 5). We also explore an alternative method to define the dissimilarity among the non-overlapping pairs in Supplementary L.
Through Stage 2, for any pair of observations, we define and calculate their dissimilarities; these dissimilarities can be organized into a dissimilarity matrix, and partitioned following Stage 1 as a block-structured dissimilarity matrix, where each block corresponding to dissimilarities within a specific pattern pair.
@@
In Stage 3, the goal is to combine the block-structured dissimilarities from Stage 2 into a global test statistic for testing $H_{0}$. Because different pattern pairs may involve varying numbers and types of shared features, they can be on incompatible scales; directly applying existing two-sample methods to the full dissimilarity matrix may therefore lead to invalid conclusions. Section 2.3 presents two aggregation strategies designed to address this issue.
After aggregation, the global statistic is compared to its null distribution using a permutation-based procedure. This is non-trivial: standard permutation preserves only the overall group sizes and ignores the distribution of missingness patterns, which can inflate type I error when the groups differ in marginal pattern probabilities. To obtain valid inference in such settings, we propose a new permutation scheme tailored to block-wise missing data. We first specify the data-generating model and then define the relevant permutation distributions.
Let $Z=(Z^{(1)},\dots,Z^{(L)})$ denote the full data vector taking values in $\prod_{l=1}^{L}\mathcal{S}^{(l)}$, and let $S\in\{0,1\}^{L}$ denote the missingness pattern. For group $g\in\{X,Y\}$, the full data follow $Z\sim F_{g}$ and the joint distribution of $(Z,S)$ is
$P_{g}(z,s)=P_{g}(S=s\mid Z=z)\,\frac{dF_{g}(z)}{dz}.$
The observed data consist of $(Z^{\mathrm{obs}},S)$, where $Z^{\mathrm{obs}}$ is the subvector of $Z$ indexed by $\{l:S^{(l)}=1\}$. All test statistics are measurable functions of $(Z^{\mathrm{obs}},S)$. Our inference is carried out under the null hypothesis $H_{0}:F_{X}=F_{Y}$, while allowing $P_{g}(S=s\mid Z=z)$ to differ between groups.

### v2

@@ Abstract
Multi-source and multi-modal datasets are increasingly common in scientific research, yet they often exhibit block-wise missingness, where entire modalities are systematically absent in some sources or no single source contains all modalities. This structured missingness poses major challenges for two-sample hypothesis testing. Standard approaches, such as imputation or complete-case analysis, may introduce bias or suffer efficiency loss, especially under missingness-not-at-random mechanisms. To address this challenge, we propose the Block-Pattern Enhanced Test, a general framework for constructing two-sample testing statistics that explicitly accounts for block-wise missingness. We show that the framework yields valid tests under a new condition allowing for missing-not-at-random mechanism. Building on this general framework, we further propose the Block-wise Rank In Similarity graph Edge-count (BRISE) test, which accommodate heterogeneous modalities using rank-based similarity graphs. Theoretically, we establish that the null distribution of BRISE converges to a $\chi^{2}$ distribution, and that the test is consistent both in the standard asymptotic regime and in the high-dimensional low-sample-size setting under mild conditions. Simulation studies demonstrate that BRISE controls the type-I error rate and achieves strong power across a wide range of alternatives. Applications to two real-world datasets with block-wise missingness further illustrate the practical utility of the proposed method.
@@ 1 Introduction
Multi-source and multi-modal datasets are increasingly common in scientific research, encompassing diverse data types collected across individuals, institutions, and time. For example, the Alzheimer’s Disease Neuroimaging Initiative (ADNI) study (Mueller et al., 2005a, b) combines multiple modalities of various types from multiple sources, including imaging, genetic, clinical data, and electronic health records (EHRs) (Johnson et al., 2023) that integrate medication histories and clinical notes to understand possible causes and progressions of Alzheimer’s Disease in different patient groups. In such scientific research, a fundamental task in statistics is to determine whether two pre-specified groups/samples come from the same distribution. Formally, let $X_{1},\ldots,X_{m}\overset{i.i.d}{\sim}F_{X}$ and $Y_{1},\ldots,Y_{n}\overset{i.i.d}{\sim}F_{Y}$ be two independent samples. The goal is to test
$H_{0}:F_{X}=F_{Y}\quad\text{versus}\quad H_{a}:F_{X}\neq F_{Y}.$
In this work, we focus on two-sample problem for multi-source and multi-modal data.
With more scientific research integrate multiple modalities from multiple sources, unsatisfactory data quality and structure are common and may bring challenges in statistical analysis, especially for two-sample testing. A key issue is block-wise missingness, where entire data modalities are absent for certain sources (Figure 1); and it is possible that no single source has all modalities available. This data structure often stems from non-uniform clinical priorities, resource constraints, or institutional protocols; and thus, the missing may not be at random but associated with the study itself. For instance, in ADNI study, site-level practices can produce systematic gaps across modalities (Xue and Qu, 2021). Similar missing patterns appear in partial or incomplete multi-view learning (Li et al., 2014; Zhang et al., 2020; Zhou et al., 2023), such as bilingual document clustering with missing translations or audiovisual tasks where only one modality is available.
@@
This challenge is especially acute in the context of two-sample testing. However, nonparametric tests for missing data are scarce; naive implementation of existing methods leads to inflated type-I error or unsatisfied power. Figure 2 summarizes three common existing strategies. Specifically, imputation method fills in missing entries using model-based predictions (Toutenburg, 1990; Zhang, 2016) but relies on strong assumptions, which can bias inference if those assumptions fail. Complete-case analysis discards incomplete samples, which reduces effective sample size (Little, 1992; Ross et al., 2020). Available-case analysis uses only variables required for a given comparison but can still suffer information loss and inefficiency (Little, 1992). Recent work has extended classical two-sample testing procedures to handle entrywise missingness, but these approaches often require missing-at-random (MAR) assumption and are conservative under block-wise missingness (Zeng et al., 2024a, b; Gretton et al., 2012; Aleksić and Milošević, 2025). Thus, a method tailored to structured block-wise mechanisms, possibly accommodating missing-not-at-random (MNAR), is still needed.
@@
Another challenge comes from integrating multiple modalities of heterogeneous data types. In the example of the ADNI study, multiple modalities are combined, including both Euclidean and non-Euclidean data. Structured data, such as clinical variables and genetic information, typically lie in Euclidean domains, whereas unstructure data, such as image data and clinical notes, may not share the same geometry as Euclidean data. As a result, a testing procedure should incorporate general and flexible distance metrics, rather than relying solely on Euclidean distances, to effectively integrate modalities of different data types in statistical analysis.
When the data are fully observed, many powerful nonparametric methods are available, including graph-based tests (Friedman and Rafsky, 1979; Schilling, 1986; Henze, 1988; Rosenbaum, 2005; Chen and Zhang, 2013; Chen and Friedman, 2017; Chen et al., 2018; Zhou and Chen, 2023), classification-based tests (Lopez-Paz and Oquab, 2017; Kim et al., 2021; Hediger et al., 2022), inter-point distance-based tests (Biswas and Ghosh, 2014; Li, 2018), and kernel-based tests (Gretton et al., 2006; Eric et al., 2007; Gretton et al., 2009, 2012; Song and Chen, 2024). Therefore, a method capable of integrating both Euclidean and non-Euclidean modalities has the potential to substantially enhance the utility of multi-source and multi-modal data.
In this work, to address structured block-wise missing mechanisms and accommodate possible MNAR missingness, we propose the Block-Pattern Enhanced Test (BPET), a general framework for two-sample testing under block-wise missing data. BPET partitions the dataset according to missingness patterns, applies pattern-aware procedures to pairs of patterns using shared observed sources, and aggregates the results into a global test statistic, allowing direct inference on partially observed data without imputation or case deletion. Importantly, BPET is not tied to any specific testing method, enabling a broad class of two-sample tests to be adapted within the framework. We show that BPET yields valid testing statistics under assumptions strictly weaker than Missing Completely At Random (MCAR), while allowing for MNAR missingness.
To better integrate modalities of heterogeneous types, we further apply the proposed framework with a rank-weighted similarity graph-based testing method (Zhou and Chen, 2023) and develop a specific test, termed the Block-wise Rank In Similarity graph Edge-count (BRISE) test. Under mild regularity conditions, we establish that the BRISE statistic converges to a $\chi^{2}$ distribution under the null hypothesis and that the test is consistent in both the standard asymptotic regime and the high-dimensional low-sample-size setting.
For finite-sample inference, we develop a tailored permutation-based procedure to approximate the null distribution. Unlike standard permutation, which shuffles all labels while fixing only total group sizes, we introduce a pattern-wise permutation that restricts permutations within each missingness pattern. We show that standard permutation can fail when groups have unequal marginal pattern distributions, whereas the proposed pattern-wise permutation yields valid inference.
In summary, our contributions are threefold: (1) We propose BPET, a general framework for two-sample testing on partially observed multi-source and multi-modal data that operates directly on block-wise missing data without imputation or case deletion; (2) Within the proposed framework, we develop the BRISE test by integrating rank-weighted similarity graph methods to accommodate modalities of heterogeneous data types; and (3) We establish the asymptotic $\chi^{2}$ null distribution and consistency of BRISE in both standard asymptotic and high-dimensional low-sample-size regimes, and propose a pattern-wise permutation scheme for valid finite-sample inference when standard permutation may fail.
@@
The remainder of the paper is organized as follows. Section 2 presents BPET and Section 3 introduces BRISE; Section 4 develops theory; Sections 5 and 6 present numerical and real data results; and Section 7 concludes.
2 Framework of Block-Pattern Enhanced Test
In this section, we introduce the general framework of BPET and a pattern-wise permutation procedure to approximate the null distribution of the test statistic. We also establish the validity of the framework under different missingness assumptions, including possible MNAR mechanisms.
2.1 General Framework
The key idea of BPET is to construct pattern-aware test statistics and aggregate them into a global statistic for two-sample testing. Since each pattern-specific statistic involves only distributions conditional on a given missingness pattern, BPET yields a valid global test as long as, under the null hypothesis, the two samples share the same conditional distribution within each pattern, allowing for possible MNAR missingness.
Stage 1: Missing-Pattern Partition.
Suppose there are $L$ sources or modalities, indexed by $l=1,\dots,L$, where each modality is defined on a domain $\mathcal{S}^{(l)}$. These domains may vary across modalities, ranging from Euclidean spaces such as $\mathbb{R}^{d_{l}}$ to non-Euclidean spaces such as networks or functional data spaces. Let $n_{P}$ denote the number of distinct missingness patterns observed in the data. Each missing pattern is represented as
@@ 2.1 General Framework
where $P_{\alpha}^{(l)}=1$ if modality $l$ is observed and 0 otherwise. For instance, when $L=3$, Figure 1 shows that $\mathcal{P}_{1}=(1,1,1)$ corresponds to complete observation; $\mathcal{P}_{2}=(1,1,0)$, $\mathcal{P}_{3}=(1,0,1)$, $\mathcal{P}_{4}=(0,1,1)$ correspond to cases with two observed modalities; and $\mathcal{P}_{5}=(1,0,0)$, $\mathcal{P}_{6}=(0,1,0)$, $\mathcal{P}_{7}=(0,0,1)$ correspond to cases with only one observed modality. In general, the number of possible nonempty patterns is at most $2^{L}-1$.
To unify notation, let $\{Z_{i}\}_{i=1}^{N}$ denote the pooled sample, where $\{Z_{i}\}_{i=1}^{m}=\{X_{j}\}_{j=1}^{m}$, $\{Z_{i}\}_{i=m+1}^{N}=\{Y_{j}\}_{j=1}^{n}$, and $N=m+n$. Throughout, we assume block-wise missingness: for each observation, variables within a modality are either fully observed or fully missing. Thus, each observation can be written as $Z_{i}=(Z_{i}^{(1)},\ldots,Z_{i}^{(L)})$, where $Z_{i}^{(l)}\in\mathcal{S}^{(l)}$ if observed, and its missingness pattern is denoted by $S_{i}=(S_{i}^{(1)},\ldots,S_{i}^{(L)})\in\{0,1\}^{L}$.
The set of observations following pattern $\mathcal{P}_{\alpha}$ is
@@
Define $\mathcal{X}^{(\alpha)}=\mathcal{Z}^{(\alpha)}\cap\{X_{i}\}_{i=1}^{m}$ and $\mathcal{Y}^{(\alpha)}=\mathcal{Z}^{(\alpha)}\cap\{Y_{i}\}_{i=1}^{n}$, so that $\mathcal{Z}^{(\alpha)}=\mathcal{X}^{(\alpha)}\cup\mathcal{Y}^{(\alpha)}$. Let $m_{\alpha}=|\mathcal{X}^{(\alpha)}|$, $n_{\alpha}=|\mathcal{Y}^{(\alpha)}|$, and $N_{\alpha}=|\mathcal{Z}^{(\alpha)}|$, where $|\cdot|$ denotes the cardinality of a set. Then
@@
We then define a dissimilarity score for any two observations from the same or different missing patterns. For observations $Z_{i}\in\mathcal{Z}^{(\alpha)}$ and $Z_{j}\in\mathcal{Z}^{(\beta)}$, if their missing patterns share at least one modality (i.e., $\mathcal{P}_{\alpha}\mathcal{P}_{\beta}^{\top}>0$), we define their dissimilarity as
@@
where $\rho_{l}(\cdot,\cdot)$ is a chosen dissimilarity measure and $\mathtt{Norm}(\cdot)$ is a normalization function. Notice that the dissimilarity between a pair of observations only uses their shared modalities. For example, in Figure 1, given $Z_{i}\in\mathcal{Z}^{(3)}$ and $Z_{j}\in\mathcal{Z}^{(5)}$ with $\mathcal{P}_{3}=(1,0,1)$ and $\mathcal{P}_{5}=(1,0,0)$, the shared modality is $\mathcal{S}^{(1)}$, and their dissimilarity reduces to $\rho(Z_{i},Z_{j})=\mathtt{Norm}\big(\rho_{1}(Z_{i}^{(1)},Z_{j}^{(1)})\big)$. If a pair of observations shares no common modalities, i.e., $\mathcal{P}_{\alpha}\mathcal{P}_{\beta}^{\top}=0$, direct computation of dissimilarity is infeasible. We adopt a simple strategy by assigning zero values to such pairs, which performs well empirically in both simulations and real data analyses (see Sections 5 and 6). An alternative strategy that further exploits indirect comparability is discussed in Supplementary M.
The dissimilarity in (1) can take various forms depending on the relationships within and across modalities. For Euclidean modalities, one may use $\rho_{l}(Z_{i}^{(l)},Z_{j}^{(l)})=\|Z_{i}^{(l)}-Z_{j}^{(l)}\|_{2}^{2}$ with $\mathtt{Norm}(x)=\sqrt{x}$, yielding Euclidean distance over shared modalities. For non-Euclidean modalities, more flexible choices of $\rho_{l}(\cdot,\cdot)$ can be adopted, such as graph diffusion distances for networks (Hammond et al., 2013), cosine similarity for embeddings (Kenter and De Rijke, 2015), or tailored measures for categorical data (Boriah et al., 2008). Different choices of $\rho_{l}$ lead to different test-specific dissimilarities and test statistics.
Through Stage 2, pairwise dissimilarities are computed and organized into a block-structured dissimilarity matrix according to missingness patterns.
@@
In Stage 3, we aggregate the block-structured dissimilarity matrix from Stage 2 into a global test statistic for testing $H_{0}$. Because different missingness patterns may involve varying sample sizes and shared modalities, direct application of standard two-sample methods to the full dissimilarity matrix may lead to invalid inference. To construct a valid test statistics, a possible strategy is to first compute pattern-wise/pattern-aware statistics based on the chosen testing method, and then aggregate information across missing patterns to construct a global test statistic. In Section 3.2, we follow this idea and propose new test statistics within the BPET framework using the RISE as the specific testing method. Other testing methods can also be adopted within the BPET framework.
2.2 A Pattern-wise Permutation Procedure
To conduct inference using the global test statistic, we approximate its null distribution via permutation. However, standard permutation preserves only the total sample sizes and ignores imbalance across missingness patterns, which can inflate type-I error under block-wise missingness. To address this issue, we propose a pattern-wise permutation scheme that permutes labels within each missingness pattern.
Let $Z=(Z^{(1)},\dots,Z^{(L)})$ denote the full data vector and $S\in\{0,1\}^{L}$ denote the missingness pattern. In two-sample testing, we have two groups/samples, i.e., Group X (samples in Group X are denoted by $X$) and Group Y (samples in Group Y are denoted by $Y$). Then for Group $G\in\{X,Y\}$, the full data follow $Z\sim F_{G}$, and the joint distribution of $(Z,S)$ is
$P_{Z,S\mid G}(z,s)=P_{S\mid Z,G}(s\mid z)\,\frac{dF_{G}(z)}{dz}.$
Under these notations, observed data consist of $(Z^{\mathrm{obs}},S)$, where $Z^{\mathrm{obs}}$ is the subvector of $Z$ indexed by $\{l:S^{(l)}=1\}$. Our inference is carried out under the null hypothesis $H_{0}:F_{X}=F_{Y}$, while allowing $P_{S\mid Z,G}(s\mid z)$ to differ between groups.